# IQ-Aware Precoding for Atomic MIMO Receivers

### v1

@@ Abstract
Leveraging the strong atom-light interaction, Rydberg atomic receivers significantly enhance the sensitivity of electromagnetic signal measurements, outperforming traditional antennas. Existing research primarily focuses on improving the architecture and signal detection algorithms of atomic receivers, while established signal processing schemes at the transmitter end have remained constant. However, these schemes fail to maximize the throughput of atomic receivers due to the nonlinearity of transmission model. To address this issue, our work represents the first attempt to design transmitter precoding in multiple-input multiple-output systems to achieve the capacity of atomic receivers. Specifically, we initially harness a strong reference approximation to convert the nonlinear magnitude-detection model of atomic receivers into a linear real-part detector. Based on this approximation, we prove that the degree of freedom is $\min\{N_{r}/2,N_{t}\}$ for a point-to-point MIMO system comprising an $N_{r}$-antenna atomic receiver and an $N_{t}$-antenna classic transmitter. To achieve the system capacity, we propose an IQ-aware fully digital precoding method. Unlike traditional complex-valued digital precoders that jointly manipulate the in-phase and quadrature (IQ) symbols, our method employs four real matrices to independently precode the IQ baseband symbols. This method is demonstrated to be optimal for atomic receivers. In addition, to eliminate the reliance on fully digital precoding architecture, we further explore IQ-aware hybrid precoding techniques. Our design incorporates a low-dimensional IQ-aware digital precoder and a high-dimensional complex analog precoder, implemented by either fully-connected or sub-connected phase shifters networks. Alternating minimization algorithms are proposed to produce the IQ-aware hybrid precoders, with the objective of closely approximating the optimal IQ-aware fully digital precoder. Simulation results validate the superiority of proposed IQ-aware precoding methods over existing techniques in atomic-receiver enabled MIMO communications.
@@
Atomic receivers, multiple-input-multiple-output (MIMO), precoding, quantum sensing.
@@ I Introduction
Ever since Hertz validated the existence of electromagnetic (EM) waves, the accurate measurement of EM waves has become the cornerstone of modern wireless communication, remote sensing, and radar [1, 2]. To date, a variety of antenna technologies have been developed, including but not limited to dipole antennas and aperture antennas. These technologies are based on a universal measurement principle, where the incident EM wave drives the antenna to generate induced currents, resulting in informative electrical signals. Although this mechanism has been used for over half a century, it imposes a notable limitation to the sensitivity (i.e., minimum detectable field strength) of traditional antennas. According to the fluctuation-dissipation theorem [3], the random movement of free electrons within metal generates indistinguishable thermal noise, known as Johnson-Nyquist noise, that contaminates electrical signals. As a result, traditional antennas typically achieve a sensitivity on the order of $1{\rm mV}/{\rm cm}/\sqrt{\rm Hz}$, making ultra-precise EM-field measurement challenging to achieve [4, 5].
@@
Recently, at the intersection of quantum sensing and wireless communication, an emerging antenna technology known as Rydberg atomic receivers has demonstrated the potential in surpassing the sensitivity limitations of traditional antennas [6, 7]. Rydberg atoms, highly excited atoms operating at high quantum states [8], possess large electric dipole moments. These atoms can strongly interact with incident EM waves, triggering electron transitions between resonant energy levels [9]. Receivers based on Rydberg atoms can monitor the strength of these transitions by leveraging quantum phenomena, such as ac-Stark shift and electromagnetically induced transparency (EIT), facilitating the detection of information-carrying EM fields [7]. This paradigm shift in EM field measurement mechanisms allows Rydberg atomic receivers to overcome the measurement limitations imposed by thermal noise, as the interaction between EM fields and atoms theoretically introduces no thermal noise. Furthermore, the quantum shot noise induced during the observation of Rydberg atoms’ quantum states is typically orders of magnitude smaller than the thermal noise of metal antennas [10]. With these advantages, atomic receivers has been reported to achieve measurement sensitivity on the order of $5{\rm nV}/{\rm cm}/\sqrt{\rm Hz}$ [11]. Thereafter, Rydberg atomic receivers show promise for use in long-range communication scenarios, such as satellite communications and space-air-ground networks [12], where detecting weak electric signals is crucial. Aligned with this vision, our work advocates the adoption of atomic receivers as a revolutionary approach to enhance wireless communication systems.
@@
Existing research on the application of Rydberg atomic receivers in detecting informative wireless signals primarily focused on the lab demonstration of single-input single-output (SISO) air interfaces. Specifically, through the interaction between EM fields and Rydberg atoms, the strength and frequency of EM waves are transduced into the intensity of electron transition, termed the Rabi frequency. By inferring symbols from Rabi frequency, researchers have successfully demonstrated the detection of amplitude-modulated (AM) and frequency-modulated (FM) signals [13, 14, 15]. To enable phase modulation, techniques such as holographic phase-sensing or quantum superheterodyne detection have been employed [16]. These techniques utilize a local oscillator (LO) to send a known reference wave that interferes with the signal wave, thereby allowing phase-modulated symbols to be extracted from the phase difference between the two waves. This method has facilitated the reception of phase-shift keying (PSK) and quadrature-amplitude modulation (QAM) symbols in communication systems [17, 18, 19], as well as the detection of micro-vibration in WiFi-sensing systems [10]. Another representative application involves the concurrent detection of multi-frequency signals spanning approximately GHz frequency bands [20, 21, 22]. This is achieved either by mixing different species of alkali atoms within a single receiver or by triggering the atoms to distinct Rydberg states.
@@
In addition to the SISO air interface, recent studies have also explored the spatial diversity and multiplexing gains of atomic receivers using Rydberg atom-based antenna arrays. In [23], a single-input multiple-output (SIMO) pipeline was established, deploying four atomic antennas to recover AM signals. A notable observation from this study was that the experimental receive signal-to-noise ratio (SNR) scales linearly with the number of atomic antennas, as expected. This atomic SIMO configuration was later implemented for angle-of-arrival estimation using the holographic phase-sensing scheme [24, 25]. To further reap the multiplexing gain, our prior work developed a multi-user signal detection model for atomic multiple-input multiple-output (MIMO) receivers [26]. We proved that atomic MIMO receivers exhibit a non-linear input-output relationship modeled by biased phase retrieval [27], in contrast to the linear transition model in classic MIMO systems. Subsequently, a maximum likelihood-based algorithm was proposed for the simultaneous detection of multi-user signals.
@@
Notably, previous efforts have primarily concentrated on improving the receiver architectures and detection algorithms, while the design of a classic transmitter capable of seamlessly collaborating with an atomic receiver remains largely uncharted. In fact, attributed to the altered input-output relationship, existing transmitter-side signal processing methods may not be optimal for atomic receivers. For instance, classic MIMO communications rely on singular value decomposition (SVD)-based precoding at the transmitter side to achieve the capacity limit $\min\{N_{r},N_{t}\}\log{\rm SNR}$ ), where $N_{r}$ and $N_{t}$ represent the number of antennas at the receiver and transmitter, respectively [28, 29]. However, this capacity is based on the linear transition model, which is not suitable for the non-linear transition model established in atomic MIMO systems. This observation raises three critical questions that remain unaddressed in existing literature:
  • What is the information-theoretical limit of an atomic MIMO receiver?
@@
  • Is there any loss in degree-of-freedom (DoF) in atomic MIMO systems due to the non-linear transition?
Addressing these questions is of pivotal significance for the seamless integration of atomic receivers into next-generation communication systems.
@@
To fill in this gap, this work represents the first attempt at analyzing and designing the transmitter precoding to approach the information-theoretic limit of Rydberg atomic MIMO systems. We consider the point-to-point MIMO communications supported by the holographic phase-sensing measurement scheme. The key contributions are summarized as follows.
  • Mutual information of atomic MIMO systems: We propose a strong-reference approximation to analyze the mutual information of atomic MIMO systems. This approximation is motivated by the observation that the reference wave is much stronger than the signal wave, as LOs are typically deployed near atomic receivers [17, 18, 19]. By discarding higher-order infinitesimal quantities, the non-linear magnitude detection model of atomic receivers can be linearized to a real-part detection model, allowing us to express the mutual information analytically. We then prove that the corresponding ergodic capacity scales as $\min\{N_{r}/2,N_{t}\}\log{\rm SNR}$ ) at high SNRs with a DoF of $\min\{N_{r}/2,N_{t}\}$. This capacity limit indicates that as long as the number of atomic antennas is at least twice the number of transmit antennas, atomic MIMO systems experience no DoF loss compared to traditional counterparts, regardless of the absence of the imaginary part information.
  • IQ-aware digital precoding: An IQ-aware digital precoding method is proposed to achieve the instantaneous capacity of atomic receivers for each channel realization. In classic MIMO systems, complex-valued digital precoders can achieve the system capacity by jointly processing the in-phase and quadrature (IQ) baseband symbols using the real and imaginary components of the precoder. However, we show that the complex-valued digital precoder is not optimal for atomic receivers due to the absence of the imaginary part of the received signal. In contrast, our IQ-aware digital precoder employs four real-valued matrices to independently precode the IQ baseband symbols before entering the IQ modulator. This offers more flexible manipulation on the distribution of radiated signals. Subsequently, by harnessing the SVD of the equivalent real-valued wireless channel, we demonstrate that an IQ-aware digital precoder achieves the instantaneous capacity of atomic receivers, with a maximal DoF of $\min\{N_{r}/2,N_{t}\}$.
@@
  • IQ-aware hybrid precoding: Despite the capacity-achieving property of IQ-aware digital precoder, it relies on a fully digital precoding architecture since analog radio-frequency (RF) circuits cannot independently process the IQ symbols. To eliminate this reliance, we further investigate the IQ-aware hybrid precoding technique, which incorporates a low-dimensional IQ-aware digital precoder realized by a baseband processor and a high-dimensional complex analog precoder realized by phase shifters (PSs). Our design focuses on two typical hybrid precoding architecture: the fully-connected (FC) and the sub-connected (SC) PS networks. We exploit the alternating minimization approach to alternatively update the digital and analog precoders, with the objective of minimizing the Frobenius distance between the IQ-aware fully digital precoder and the designed IQ-aware hybrid precoder. Numerical results validate that the FC IQ-aware hybrid precoder can closely approach the optimal IQ-aware digital precoder. Moreover, we show that the proposed IQ-aware hybrid precoding schemes significantly outperform classical hybrid precoding methods in atomic-receiver based MIMO systems, demonstrating the effectiveness of the proposed designs.
@@
Organization: The remainder of this paper is organized as follows. The system model of atomic MIMO receivers is provided in Section II. In Section III, we analyze the mutual information of atomic receivers and the limit of ergodic capacity. The IQ-aware fully digital precoding for achieving the instantaneous capacity is discussed Section IV. The IQ-aware hybrid precoding for the FC and SC architectures are elaborated in Section V. In Section VI, we validate the proposed design by numerical results. Finally, conclusions are drawn in Section VII.
@@ II Review of Rydberg Atomic MIMO System
Figure 1: Layout of the Rydberg atomic MIMO system.
We consider an atomic MIMO communication system, as illustrated in Fig. 1. The transmitter employs a conventional $N_{t}$-element antenna array for data transmission. The time-domain radiated signal from the $n$-th antenna, $n\in\{1,2,\cdots,N_{t}\}$, is represented as $x_{n}(t)=A_{n}\cos(\omega t+\gamma_{n})$. To realize phase demodulation at the receiver side, we adopt the holographic-phase sensing measurement, where a LO is introduced to radiate predetermined reference signals to the atomic receiver. Let the reference signal, $x_{R}(t)$, share the same frequency with the transmitted signal, and thereby it is expressed as $x_{R}(t)=A_{R}\cos(\omega t+\gamma_{R})$.
The atomic receiver deploys $N_{r}$ Rydberg atomic antennas for detecting the wireless signals. Each atomic antenna consists of a vapor cell filled with Alkali atoms, followed by a photodetector (PD). These atoms are excited to Rydberg states for sensing the incident EM wave via atom-light interaction, while the sensing results are read out by PDs [30]. Adopting the multi-path wireless channel model, the EM wave, ${\mathbf{E}}_{m}(t)$, impinging on the $m$-th atomic antenna, $m\in\{1,2,\cdots,N_{r}\}$, is written as
@@
Here, $L$ denotes the number of paths between the transmitter and receiver. Variables $\bm{\epsilon}_{mnl}$, $\rho_{mnl}$, and $\varphi_{mnl}$ represent the polarization direction, path loss, and phase shift of the propagation from the $n$-th transmit antenna to the $m$-th atomic antenna via the $\ell$-th path. Besides, variables $\bm{\epsilon}_{R,m}$, $\rho_{R,m}$, and $\varphi_{R,m}$ refer to the channel parameters pertaining to the reference signal. Since the LO is typically deployed nearby the receiver, we directly employ the line-of-sight path to model the LO-to-receiver channel. Suppose the signal frequency of the incident EM wave, $\omega$, is tuned resonant with the transition frequency between the ground and excited energy levels of Rydberg atoms. Governed by the time-dependent Schrödinger equation [31], the electrons will oscillate between the ground and excited energy levels, a phenomenon referred to as electron transition. The intensity of electron transition is determined by the Rabi frequency [31], $\Omega_{m}$, given as
@@
where $\bm{\mu}_{\rm RF}$ represents the electric dipole moment between Rydberg states, and $\hbar$ denotes the reduced Planck constant. Applying the rotating wave approximation [31], our prior work [26] proved that the vector ${\mathbf{E}}_{{\rm RF},m}$ can be equivalently expressed as the down-converted complex form of ${\mathbf{E}}_{m}(t)$, i.e.,
@@
where the frequency, $\omega$, is cancelled due to electron transition. Given that the wireless signals are embedded into Rabi frequencies, $\{\Omega_{m}\}$, accurate signal detection hinges on measuring their values using practical implementations, such as the EIT effects. Specifically, the electron transition of Rydberg atoms changes the susceptibility of the vapor cell [7]. By steering a probe laser to penetrate through the vapor cell and observing the dynamic change of its spectrum using a PD, the Rabi frequency can be read out. We denote the measurement of $[\Omega_{1},\cdots,\Omega_{N_{r}}]^{T}$ as a vector, ${\mathbf{y}}=[y_{1},\cdots,y_{N_{r}}]^{T}$, followed by
@@
Here, ${\mathbf{H}}\in\mathbb{C}^{N_{r}\times N_{t}}$ represents the effective channel, with the $(m,n)$-th entry given as ${\mathbf{H}}_{m,n}=\sum_{\ell=1}^{L}\frac{1}{\hbar}\bm{\mu}_{\rm RF}^{T}\bm{\epsilon}_{mnl}\rho_{mnl}e^{j\varphi_{mnl}}$, the vector ${\mathbf{x}}=[x_{1},x_{2},\cdots,x_{N_{t}}]^{T}$ with $x_{n}=A_{n}e^{j\gamma_{n}}$ denotes the transmitted signal in the complex form, and ${\mathbf{r}}=[r_{1},r_{2},\cdots,r_{N_{r}}]^{T}$ with $r_{m}=\frac{1}{\hbar}\bm{\mu}_{\rm RF}^{T}\bm{\epsilon}_{R,m}\rho_{R,m}A_{R}e^{j(\varphi_{R,m}+\gamma_{R})}$, denotes the received reference signal.
Finally, by invoking the law of large numbers, the measurement noise can be modelled as the circularly symmetric Gaussian distribution, ${\mathbf{w}}\sim\mathcal{CN}(0,\sigma^{2}{\mathbf{I}}_{N_{r}})$. As a result, we arrive at the following input-output relationship for atomic MIMO systems:
@@
which is a non-linear magnitude-detection model. To analyze the achievable capacity, $C$, of Rydberg receivers, this article focuses on the transmitter design to maximize the mutual information between ${\mathbf{y}}$ and ${\mathbf{x}}$:

### v2

# MIMO Precoding for Rydberg Atomic Receivers

@@ Abstract
Leveraging the strong atom-light interaction, a Rydberg atomic receiver can measure radio waves with extreme sensitivity. Existing research primarily focuses on improving the architecture and signal detection capability of atomic receivers, while traditional signal processing schemes at the transmitter side have remained unchanged. As a result, these schemes fail to maximize the throughput of atomic receivers, given that the coupling between atomic dipole moment and radio-wave magnitude results in a nonlinear transmission model in contrast to the traditional linear one. To address this issue, our work proposes to design customized precoding techniques for atomic multiple-input-multiple-output (MIMO) systems to achieve the channel capacity. A strong-reference approximation is initially proposed to linearize the nonlinear transition model of atomic receivers. This facilitates the derivation of atomic-MIMO channel capacity as $\min(N_{r}/2,N_{t})\log({\rm SNR})$ at high signal-to-noise ratios (SNRs) for $N_{r}$ receive atomic antennas and $N_{t}$ classic transmit antennas. Then, a new digital precoding technique, termed In-phase-and-Quadrature (IQ) aware precoding is presented, which features independent processing of I/Q data streams using four real-valued matrices. The design is shown to be capacity-achieving for the atomic MIMO system. In addition, for the case of large-scale MIMO system, we extend the preceding fully-digital precoding design to the popular hybrid precoding architecture, which cascades a classical analog precoder with a low-dimensional version of the proposed IQ-aware digital precoder. By alternatively optimizing the digital and analog parts, the hybrid design is able to approach the performance of the optimal IQ-aware fully digital precoding. Simulation results validate the superiority of proposed IQ-aware precoding methods over existing techniques in the context of atomic MIMO communication.
@@
Atomic receivers, multiple-input-multiple-output (MIMO), precoding.
@@ I Introduction
Ever since Hertz validated the existence of electromagnetic (EM) waves, the accurate measurement of EM waves has become the cornerstone of modern wireless communication, remote sensing, and radar [1, 2]. To date, a variety of antenna technologies have been developed, e.g., dipole antennas and aperture antennas. These technologies are based on the same measurement principle: an incident EM wave drives an antenna to generate induced currents, which results in informative electrical signals. This mechanism imposes a notable limitation on the sensitivity (i.e., minimum detectable field strength) of traditional antennas. According to the fluctuation-dissipation theorem [3], the random movement of free electrons within metal generates indistinguishable thermal noise, known as Johnson-Nyquist noise, that contaminates electrical signals. As a result, traditional antennas typically achieve a sensitivity on the order of $1{\rm uV}/{\rm m}/\sqrt{\rm Hz}$ [4], making ultra-precise EM-field measurement challenging to achieve [5, 6].
@@
Recently, at the intersection of quantum sensing and wireless communication, an emerging antenna technology known as Rydberg atomic receiver has shown the potential in surpassing the sensitivity limitations of traditional antennas [7, 8]. Rydberg atoms, referring to excited atoms operating at high quantum states [9], can strongly interact with incident EM waves, triggering electron transitions between resonant energy levels [10]. Receivers based on Rydberg atoms can monitor the strength of these transitions by leveraging quantum phenomena, such as ac-Stark shift and electromagnetically induced transparency (EIT), to detect information carried by the fields [8]. Rydberg atomic receivers can overcome the limitations imposed by thermal noise, as the interaction between EM fields and atoms theoretically introduces no thermal noise. Furthermore, the quantum shot noise induced during the observation of Rydberg atoms’ quantum states is typically orders of magnitude smaller than the thermal noise of metal antennas [5]. With these advantages, atomic receivers can theoretically achieve measurement sensitivity determined by standard quantum limit, e.g., $1{\rm nV}/{\rm m}/\sqrt{\rm Hz}$ [11]. Thus, such receivers show a promise for use in long-range communication scenarios, such as satellite communications and space-air-ground integrated networks [12], where detecting weak signals is crucial. Aligned with this vision, our work advocates the adoption of atomic receivers as a revolutionary approach to enhance wireless communication systems.
@@
Existing research on the application of Rydberg atomic receivers in detecting modulated wireless signals primarily focuses on the lab demonstration of single-input-single-output (SISO) air interfaces. Specifically, the EM field is captured by an atomic parameter called Rabi frequency, which characterizes the electron transition intensity. By inferring symbols from Rabi frequencies, researchers have successfully demonstrated the detection of amplitude-modulated (AM) and frequency-modulated (FM) signals [13, 14, 15, 16]. To enable phase detection, techniques such as holographic phase-sensing or heterodyne measurement have been employed [17]. These techniques utilize a local oscillator (LO) to send a known reference wave that interferes with the signal wave, thereby allowing phase-modulated symbols to be extracted from the phase difference between the two waves. This method has facilitated the implementation of phase-shift keying (PSK) and quadrature-amplitude modulation (QAM) communication systems [18, 19, 20]. Another representative application involves the concurrent detection of multi-frequency signals spanning GHz frequency bands [21, 22, 23]. This is achieved either by mixing different species of alkali atoms within a single receiver or by exciting the atoms to distinct Rydberg quantum states.
@@
In addition to the SISO air interface, recent studies have also explored the spatial diversity and multiplexing gains of atomic antenna arrays. In [24], a single-input-multiple-output (SIMO) system was designed with four atomic antennas to recover AM signals. A notable observation from this study was that the experimental receive signal-to-noise ratio (SNR) scales linearly with the number of atomic antennas, as in traditional systems. This atomic SIMO configuration was later implemented for angle-of-arrival estimation using the heterodyne measurement scheme [25, 26]. To further reap the multiplexing gain, our prior work developed a multi-user signal detection model for atomic multiple-input-multiple-output (MIMO) receivers [27]. We proved that atomic MIMO receivers exhibit a non-linear input-output relationship modeled by biased phase retrieval [28], in contrast to the linear transition model in classic MIMO systems. Subsequently, a maximum likelihood-based algorithm was proposed for the simultaneous detection of multi-user signals.
@@
Previous efforts have primarily concentrated on improving the receiver architectures and detection algorithms, while the design of a classic transmitter optimized for an atomic receiver remains largely uncharted. Attributed to the altered input-output relationship, existing transmitter-side signal processing methods may not be optimal for atomic receivers. For instance, classic MIMO communications rely on singular value decomposition (SVD)-based precoding at the transmitter side to achieve the channel capacity $\min(N_{r},N_{t})\log{\rm SNR}+\mathcal{O}(1)$ ), where $N_{r}$ and $N_{t}$ represent the numbers of antennas at the receiver and transmitter, respectively [29, 30]. However, this capacity is based on the linear transition model, while the atomic counterpart is non-linear. This difference raises three critical questions that remain unaddressed in the existing literature:
  • What is the information-theoretical limit of an atomic MIMO system?
@@
  • Is there any loss in degree-of-freedoms (DoFs) in atomic MIMO systems due to the non-linear transition?
@@
Addressing these questions is pivotal for unleashing the full potential of atomic receivers for next-generation communication systems. To this end, this work represents the first attempt on designing the transmit precoding to approach the channel capacity of Rydberg atomic MIMO systems. We consider the point-to-point MIMO communications supported by the mentioned heterodyne measurement scheme [17]. The key contributions are summarized as follows.
  • Channel capacity with atomic MIMO receivers: The non-linear input-output relationship of atomic MIMO receivers poses a challenge to express its mutual information analytically. To overcome this challenge, we propose a strong-reference approximation to linearize the transition model, which is motivated by the fact that the reference wave is much stronger than the signal wave due to the proximity of the LOs to atomic receivers [18, 19, 20]. Based on this approximation, we can convert the non-linear magnitude detection model of atomic receivers into a linear real-part detection model by discarding high-order infinitesimal quantities, which allows us to obtain the expression of channel capacity.
  • IQ-aware digital precoding: An IQ-aware digital precoding is proposed to achieve the channel capacity. Traditional MIMO precoding widely adopts a complex-valued precoder to process in-phase and quadrature (IQ) baseband symbols. This method can only produce circularly-symmetric Gaussian distributed signals, which are not optimal for the magnitude-detection based atomic MIMO receivers. In contrast, the proposed IQ-aware digital precoding employs four real-valued matrices to independently process the I/Q baseband symbols, which has the ability to produce general Gaussian distributed signals. The design is shown to achieve the channel capacity with atomic receivers, which is proven to be $\min(N_{r}/2,N_{t})\log{\rm SNR}$ ) at high SNRs. This capacity result reveals that as long as the number of receive atomic antennas, $N_{r}$, is at least twice the number of transmit traditional antennas, $N_{t}$, atomic MIMO receivers experience no loss in spatial DoFs with respect to the traditional counterparts.
@@
  • IQ-aware hybrid precoding: For the case of large-scale MIMO systems, we extend the preceding fully digital precoding design to the popular hybrid precoding architecture [31]. Our design cascades the high-dimensional complex analog precoder implemented using phase shifters (PSs) with a low-dimensional version of the proposed IQ-aware digital precoder, where both the fully-connected (FC) and the sub-connected (SC) PS networks are taken into account. We exploit the alternating minimization approach to alternatively update the digital and analog precoders, with the objective of minimizing the Frobenius distance between the previous IQ-aware fully digital precoder and the current IQ-aware hybrid precoder. Numerical results validate their close performance. Moreover, we show that the proposed IQ-aware hybrid precoding schemes significantly outperform classical hybrid precoding methods in the context of atomic MIMO systems.
@@
Organization: The remainder of this paper is organized as follows. The system model of atomic MIMO receivers is provided in Section II. In Section III, we formulate the system channel capacity using strong-reference approximation. The IQ-aware digital precoding for achieving the channel capacity is discussed Section IV. The IQ-aware hybrid precoding design for the FC and SC architectures are presented in Section V. In Section VI, numerical results are presented, followed by conclusions drawn in Section VII.
@@ II Review of Rydberg Atomic MIMO System
Figure 1: Rydberg atomic MIMO system.
To begin with, we briefly review the system model of an atomic MIMO communication system. Readers can refer to our prior work [27] for a more comprehensive understanding of Rydberg atomic receivers. As illustrated in Fig. 1, the transmitter employs a conventional $N_{t}$-element antenna array for data transmission. The time-domain radiated signal from the $n$-th antenna, $n\in\{1,2,\cdots,N_{t}\}$, is represented as $x_{n}(t)=A_{n}\cos(\omega t+\gamma_{n})$. To realize phase demodulation at the receiver side, the heterodyne measurement scheme, also known as the holographic-phase sensing measurement, is adopted [17]. This scheme introduces a LO as an auxiliary source to radiate predetermined reference signals to the atomic receiver. For ease of discussion, we assume that the reference signal, $x_{R}(t)$, shares the same frequency with the transmitted signal, and thereby it is expressed as $x_{R}(t)=A_{R}\cos(\omega t+\gamma_{R})$.
The atomic receiver deploys $N_{r}$ Rydberg atomic antennas for detecting the incident signals. Each antenna consists of a vapor cell filled with Alkali atoms. These atoms are excited to Rydberg states via two-photon transition process using a probe laser and a coupling laser[5]. The incident EM wave can strongly interact with the electric dipole moments of these Rydberg atoms, while the interaction results are read out by photodetectors (PDs) to perform signal detection [32]. Mathematically speaking, by adopting the multi-path wireless channel model, the EM wave ${\mathbf{E}}_{m}(t)$ impinging on the $m$-th atomic antenna, $m\in\{1,2,\cdots,N_{r}\}$, is written as
@@
Here, $L$ denotes the number of transmitter-to-receiver channel paths. Variables $\bm{\epsilon}_{mnl}$, $\rho_{mnl}$, and $\varphi_{mnl}$ represent the polarization direction, path loss, and phase shift of the radio wave propagating from the $n$-th transmit antenna to the $m$-th atomic antenna via the $\ell$-th path. Variables $\bm{\epsilon}_{R,m}$, $\rho_{R,m}$, and $\varphi_{R,m}$ refer to the channel parameters pertaining to the reference signal. Since the LO is typically deployed nearby the receiver, we directly employ the line-of-sight path to model the LO-to-receiver channel.
Suppose the signal frequency of the incident EM wave, $\omega$, is tuned to resonant with the transition frequency, $\omega_{ge}$, between two specific energy levels of Rydberg atoms, denoted as |$\ket{g}$ and |$\ket{e}$. Then, governed by the Lindblad master equation (Schrödinger equation in open quantum systems) [33], the electrons of Rydberg atoms will oscillate between these two energy levels, a phenomenon referred to as electron transition [27]. The transition intensity is determined by a physical parameter called Rabi frequency [33]:
@@
where $\bm{\mu}_{\rm RF}$ represents the electric dipole moment of Rydberg atoms, and $\hbar$ denotes the reduced Planck constant. Applying the rotating wave approximation [33], our prior work [27] reveals that the vector ${\mathbf{E}}_{{\rm RF},m}$ can be equivalently expressed as the down-converted complex form of ${\mathbf{E}}_{m}(t)$, i.e.,
@@
where the signal frequency, $\omega$, is physically cancelled by its resonant transition frequency, $\omega_{ge}=\omega$. Given that the signals, $\{{\mathbf{E}}_{{\rm RF},m}\}$, are embedded into Rabi frequencies, $\{\Omega_{m}\}$, efficient signal detection hinges on measuring $\{\Omega_{m}\}$ accurately. This is popularly accomplished by a quantum coherence process called electromagnetically induced transparency (EIT) spectroscopy [8]. Briefly speaking, the electron transition process changes the rate at which Rydberg atoms absorb photons from an incoming laser [8]. Thereby, by steering a probe laser to penetrate through the vapor cell and observing the dynamic change of its spectrum using a PD, the Rabi frequency can be read out. We denote the measurement of $[\Omega_{1},\cdots,\Omega_{N_{r}}]^{T}$ as a vector, ${\mathbf{y}}=[y_{1},\cdots,y_{N_{r}}]^{T}$, followed by
@@
Here, ${\mathbf{H}}\in\mathbb{C}^{N_{r}\times N_{t}}$ is the effective channel matrix, whose $(m,n)$-th entry is ${\mathbf{H}}_{m,n}=\sum_{\ell=1}^{L}\frac{1}{\hbar}\bm{\mu}_{\rm RF}^{T}\bm{\epsilon}_{mnl}\rho_{mnl}e^{j\varphi_{mnl}}$, the vector ${\mathbf{x}}=[x_{1},x_{2},\cdots,x_{N_{t}}]^{T}$ with $x_{n}=A_{n}e^{j\gamma_{n}}$ denotes the transmitted signal in the complex form, and ${\mathbf{r}}=[r_{1},r_{2},\cdots,r_{N_{r}}]^{T}$ with $r_{m}=\frac{1}{\hbar}\bm{\mu}_{\rm RF}^{T}\bm{\epsilon}_{R,m}\rho_{R,m}A_{R}e^{j(\varphi_{R,m}+\gamma_{R})}$, denotes the received reference signal.
Finally, by invoking the law of large numbers, the measurement noise can be modelled as the complex Gaussian distribution, ${\mathbf{w}}\sim\mathcal{CN}(0,\sigma^{2}{\mathbf{I}}_{N_{r}})$. As a result, we arrive at the following input-output relationship for atomic MIMO systems:
@@
which is a non-linear magnitude-detection model. To analyze its achievable capacity, $C$, this article focuses on the transmitter precoding design to maximize the mutual information between ${\mathbf{y}}$ and ${\mathbf{x}}$: